# Efficient Bilinear Attention-based Fusion for Medical Visual Question Answering

### v1

@@ Abstract
Medical Visual Question Answering (MedVQA) has gained increasing attention at the intersection of computer vision and natural language processing. Its capability to interpret radiological images and deliver precise answers to clinical inquiries positions MedVQA as a valuable tool for supporting diagnostic decision-making for physicians and alleviating the workload on radiologists. While recent approaches focus on using unified pre-trained large models for multi-modal fusion like cross-modal Transformers, research on more efficient fusion methods remains relatively scarce within this discipline. In this paper, we introduce a novel fusion model that integrates Orthogonality loss, Multi-head attention and Bilinear Attention Network (OMniBAN) to achieve high computational efficiency and strong performance without the need for pre-training. We conduct comprehensive experiments and clarify aspects of how to enhance bilinear attention fusion to achieve performance comparable to that of large models. Experimental results show that OMniBAN outperforms traditional models on key MedVQA benchmarks while maintaining a lower computational cost, which indicates its potential for efficient clinical application in radiology and pathology image question answering.
@@ I Introduction
Medical Visual Question Answering (MedVQA) is an emerging field within multi-modal artificial intelligence that adapts the principles of general Visual Question Answering (VQA) to meet the specific demands of the medical domain. The primary goal of MedVQA is to support healthcare professionals by automatically generating accurate answers to clinical questions based on medical images, thereby assisting in clinical decision-making and relieving workload. This task involves the fusion of computer vision and natural language processing techniques to analyze visual data alongside natural language questions, enabling contextually relevant and clinically accurate responses.
MedVQA presents unique challenges due to the specialized nature of medical images and the technical terminology prevalent in clinical inquiries. Effectively addressing these complexities requires models capable of capturing fine-grained details and sophisticated relationships between visual and textual data. While traditional VQA models such as VGGNet [1], ResNet [2], GRU [3], and LSTM [4] have been adapted for MedVQA tasks, they often fall short in the medical context due to limited labeled data and the need for robust generalization across diverse clinical scenarios.
Despite the recent success of multimodal fusion techniques in enhancing MedVQA performance, there is a significant gap in research focusing on computationally efficient fusion methods. Transformer-based models, especially cross-modal Transformers, have demonstrated strong fusion capabilities and have been widely adopted in this domain. However, these large unified models come with substantial computational demands, making them less suitable for real-time clinical applications where computational efficiency is essential. This motivates the need for exploring alternative fusion techniques that maintain high performance while reducing computational complexity, particularly for applications constrained by processing power.
In this paper, we introduce a novel model, OmniBANet, that integrates Orthogonality loss, Multi-head attention, and a Bilinear Attention Network (OMniBAN). Our model is specifically designed to address the challenges of MedVQA by achieving a high level of computational efficiency and robust performance without requiring pre-training, making it suitable for clinical scenarios with resource constraints. Through comprehensive experiments, we demonstrate that OmniBANet achieves competitive results on key MedVQA benchmarks while operating at a lower computational cost than Transformer-based models. This balance of efficiency and accuracy indicates OmniBANet’s potential for effective application in medical image question answering across radiology and pathology domains.
@@ II Related Work
Figure 1: Overview of our proposed Orthogonal Multi-head Bilinear Attention Network (OMniBAN).
@@ II-A Medical Visual Question Answering
Medical Visual Question Answering (MedVQA) is an emerging research area within multimodal artificial intelligence that applies the general principles of Visual Question Answering (VQA) to the medical domain. This field combines computer vision and natural language processing techniques to analyze and understand medical images in conjunction with natural language questions, with the goal of generating accurate answers that can assist in clinical decision-making.
Initial research efforts in MedVQA adopt VQA models that have proven effective in general VQA, and adapt them for medical applications.
In terms of image feature extraction, researchers often rely on pre-trained models like VGGNet [1] and ResNet [2], which are fine-tuned for the specific task of MedVQA. On the text side, methods like GRU [3] and LSTM [4] are commonly used to extract textual features, while some approaches incorporate additional semantic information derived from medical corpora to enhance the embeddings used for question representation.
@@ II-B Multi-modal Fusion
In the Visual Question Answering task, multi-modal fusion plays a crucial role by integrating visual and textual features to enable accurate classification. The performance of VQA models largely depends on the effective extraction of features within modalities and the subsequent fusion across modalities.
Recent approaches to multi-modal fusion in VQA have introduced various methods to enhance the integration of visual and textual features. A foundational method is SAN[7], which iteratively refines attention on relevant image regions based on the given question. Bilinear pooling has been a major focus in improving fusion by capturing complex interactions between modalities. Yu et al. [8] proposed MFB to address the high computational cost of bilinear pooling by factorizing the interaction into two low-rank matrices, preserving performance while reducing complexity. MUTAN[9] builds on this by applying Tucker decomposition to further compress the bilinear tensor, resulting in a more compact and efficient multi-modal representation. To make full use of bilinear attention maps, Kim et al.[10] proposed BAN, which capture intricate dependencies between visual and textual features efficiently.
The advent of Transformer[11] has brought a significant shift in multi-modal fusion strategies. Methods like LXMERT[12], hi-VQA[13] and MCAN[14] utilize cross-modality Transformers with separate encoders for vision and language, followed by a cross-modality encoder to fuse the extracted features. In the domain of Medical Visual Question Answering, Liu et al.[15] employs a Transformer-based architecture that directly fuse image and text features to generate joint representation. While Transformers are effective for fusion, they require more computational resources and are more complex compared to bilinear pooling-based methods. This makes balancing performance and computational cost particularly important in Medical Visual Question Answering, where medical data is often limited.
@@ III-B Multi-modal Feature Extraction
Question Encoder. We use the pre-trained BioBERT model [18] as our question encoder due to its suitability for processing complex biomedical language. BioBERT, based on the BERT architecture, has been fine-tuned on a large biomedical corpus, enabling it to capture the nuanced and specialized terminology that characterizes the biomedical domain. This adaptation allows BioBERT to generate more accurate representations of medical language than general-purpose language models.
@@
While BiomedCLIP also offers a text encoder, we chose not to use it for question encoding in this work. Although the BiomedCLIP Text Encoder aligns image and text features within a shared feature space, it lacks the word-level granularity necessary for capturing detailed linguistic information. This level of detail is critical for the fusion methods introduced in Section III-C, which benefit from precise, word-level representations in the textual data.
@@ III-C Orthogonal Multi-head Bilinear Attention Network
Our proposed Orthogonal Multi-head Bilinear Attention Network (OMniBAN) integrates a single-layer multi-head self-attention mechanism with bilinear attention networks to efficiently fuse visual and textual features for Medical Visual Question Answering. This design enables the model to capture complex intra-modal and cross-modal interactions effectively while maintaining computational efficiency. By leveraging orthogonal multi-head attention, OMniBAN enhances feature diversity and maximizes information extraction across modalities without requiring extensive pre-training.
@@ III-C2 Cross-modal Bilinear Attention
where $\circ$ denotes the Hadamard (element-wise) product, allowing for fine-grained alignment between corresponding elements in the image and question features.
@@
In this formulation, each attention glimpse $h$ learns specialized cross-modal relationships, allowing the model to capture diverse interaction patterns between visual and textual features.
@@ III-C3 Orthogonality Loss
To ensure diverse information captured by the model, we introduce an Orthogonality Loss[19] to encourage each attention glimpse to focus on unique aspects of the input. This loss reduces redundancy across glimpses, which is particularly beneficial for the complex nature of Medical Visual Question Answering (MedVQA), where both image and text data contain rich, detailed information. Additionally, it helps address with over-fitting, which is common in the field of MedVQA.
In this work, Orthogonality Loss is applied to the attention distributions obtained from the bilinear attention mechanism, as shown in Algorithm 1. For each pair of attention distributions (glimpses), the inner product of their normalized vectors is computed and squared, with these values summed to form the final orthogonality loss. This approach promotes orthogonality between attention glimpses, reducing overlap and increasing diversity in the attended features across different heads.
@@
10:$\mathcal{L}_{\text{ortho}}\leftarrow 0$
@@
14:     Update orthogonality loss: $\mathcal{L}_{\text{ortho}}\leftarrow\mathcal{L}_{\text{ortho}}+\rho^{2}$
@@ III-C4 Classifier and Prediction
The joint representation output is then fed into a classifier to predict the most probable answer. A simple feed-forward neural network is used as the classifier, comprising two fully connected layers with an intermediate activation function. This prediction can be represented as:
$\hat{a}=\text{Classify}(\mathbf{h}^{\text{joint}})$ (9)
@@
The main loss function for this task is Binary Cross-Entropy with Logits Loss, a common choice for multi-label classification. The total loss used during training combines the main classification loss with the Orthogonality Loss, which encourages both accurate predictions and diversified attention features.
$L=L_{class}+\alpha\cdot L_{ortho}$ (10)
where $\alpha$ is the threshold for Orthogonality Loss.
III-C5 Theoretical Analysis of Computational Complexity
In terms of computational complexity, both the OMniBAN model and pure cross-modal Transformer-based models compute attention across all pairs of elements in both modalities, but they differ in how they handle these interactions and in the subsequent processing steps.
In a typical cross-modal Transformer-based model, the attention mechanism has a complexity of $O(N_{v}\times N_{q}\times d)$ per layer, where $N_{v}$ and $N_{q}$ are the number of visual and textual elements, respectively, and $d$ is the feature dimension. This complexity arises because the Transformer computes attention weights for every possible pair of elements in the two modalities, followed by a weighted sum that produces a new set of feature representations.
In contrast, the Bilinear Attention Network used in OMniBAN also calculates interactions across all pairs, but it does so in bilinear space by factorizing the interaction into lower-dimensional spaces. Specifically, BAN employs learnable projection matrices $\mathbf{W_{v}}$ and $\mathbf{W_{q}}$, which project the image and text features into a joint space where the interactions are computed. While this operation still involves all pairs of elements, the subsequent bilinear pooling and the use of lower-dimensional projections (i.e., $d_{m}$ instead of $d$) make the computation more efficient in practice. The complexity of this bilinear interaction is $O(N_{v}\times N_{q}\times d_{m})$, where $d_{m}$ is typically smaller than $d$, particularly in cases where the dimensionality reduction is significant.
Additionally, in Transformer-based models, after the attention computation, the features are passed through multiple layers of feed-forward networks, which further add to the computational burden. In contrast, OMniBAN uses only a single layer of multi-head self-attention and leverages bilinear attention’s direct aggregation of multimodal features. It can achieve performance comparable to Transformer-based models without the need for extensive post-attention processing or pre-training.
Therefore, while both models compute interactions across all pairs, OMniBAN achieves a lower overall complexity by reducing the dimensionality of the interaction space and simplifying the subsequent processing steps.
@@ IV Experiments
IV-A Dataset and Metric
We use VQA-RAD dataset in our experiments, which contains 3,515 QA pairs based on 315 radiology images. The questions in this dataset are categorized into two types: Closed and Open. Closed questions are those with a limited set of possible answers, most commonly yes/no, while Open questions are more varied and do not restrict the type of response. The dataset is split into a training set with 3,064 question–answer pairs and a test set with 451 pairs.
For the evaluation metric, we primarily use accuracy, as MedVQA can be viewed as a multi-class classification task. Accuracy measures the proportion of correctly predicted answers out of the total number of questions, providing a straightforward and widely accepted indicator of model performance in this context.

### v3

@@ Abstract
Medical Visual Question Answering (MedVQA) has attracted growing interest at the intersection of medical image understanding and natural language processing for clinical applications. By interpreting medical images and providing precise answers to relevant clinical inquiries, MedVQA has the potential to support diagnostic decision-making and reduce workload across various fields like radiology. While recent approaches rely heavily on unified large pre-trained Visual-Language Models, research on more efficient fusion mechanisms remains relatively limited in this domain. In this paper, we introduce a fusion model, OMniBAN, that integrates Orthogonality loss, Multi-head attention, and a Bilinear Attention Network to achieve high computational efficiency as well as solid performance. We conduct comprehensive experiments and demonstrate how bilinear attention fusion can approximate the performance of larger fusion models like cross-modal Transformer. Our results show that OMniBAN requires fewer parameters (approximately 2/3 of Transformer-based Co-Attention) and substantially lower FLOPs (approximately 1/4), while achieving comparable overall performance and even slight improvements on closed-ended questions on two key MedVQA benchmarks. This balance between efficiency and accuracy suggests that OMniBAN could be a viable option for real-world medical image question answering, where computational resources are often constrained.
@@ I Introduction
Medical Visual Question Answering (MedVQA) is an emerging field within multi-modal artificial intelligence that adapts the principles of general Visual Question Answering (VQA) to meet the specific demands of the medical domain. The primary goal of MedVQA is to support healthcare professionals by automatically generating accurate answers to clinical questions based on medical images, thereby assisting in clinical decision-making and relieving workload. This task involves the fusion of computer vision and natural language processing techniques to analyze visual data alongside natural language questions to enable contextually relevant and clinically accurate responses.
Compared to general domain Visual Question Answering, MedVQA presents unique and significant challenges. At the image level, medical images such as those in radiology or pathology often exhibit subtle differences within small regions, where even minor pixel variations can represent completely different diagnostic findings (e.g., tiny lesions). At the text level, clinical inquiries are characterized by specialized terminology and complex language structures. This necessitates language models equipped with domain-specific knowledge to accurately interpret the semantics of medical questions. Given these challenges, the multi-modal fusion module is important, as it must ensure the effective integration of information from these distinct modalities without losing key information.
Despite the recent success of multimodal fusion techniques in enhancing MedVQA performance, there is a significant gap in research focusing on computationally efficient fusion methods. Transformer-based models, especially cross-modal Transformers, have demonstrated strong fusion capabilities and have been widely adopted in this domain. However, these large unified models come with substantial computational demands, making them less suitable for real-time clinical applications or small MedVQA model training, where computational efficiency is essential. Also, a common training strategy for MedVQA involves directly utilizing the embeddings from frozen visual and textual encoders, therefore, the primary computational cost of training concentrates on the fusion network. This motivates the need for exploring alternative fusion techniques that maintain high performance while reducing computational complexity.
In this paper, we propose an efficient fusion framework called OMniBAN, which combines Orthogonality loss, Multi-head attention, and a Bilinear Attention Network. Our approach is designed to deliver comparable performance on MedVQA tasks at a lower training and computational cost. Through extensive experiments, we show that OMniBAN achieves similar results compared with large Transformer-based fusion models on key MedVQA benchmarks, yet it requires fewer computational resources. This efficiency–accuracy trade-off positions OMniBAN as a promising solution for real-world medical image question answering, particularly in radiological Visual Question Answering.
@@ II Related Work
Figure 1: Overview of our proposed Orthogonal Multi-head Bilinear Attention Network (OMniBAN). The frozen visual and textual backbones encode the input image and question independently, and then the core OMniBAN fusion module fuses these features using multi-head attention for within-modality refinement and bilinear attention to capture cross-modal interactions. Orthogonality loss is adopted to encourage diverse attention patterns among glimpses during training.
@@ II-A Medical Visual Question Answering
Medical Visual Question Answering (MedVQA) is an emerging research area within multimodal artificial intelligence that applies the general principles of Visual Question Answering (VQA) to the medical domain. This field combines medical image understanding and natural language processing techniques to analyze and understand medical images in conjunction with natural language questions, with the goal of generating accurate answers useful for clinical decision-making and diagnostic assistance.
Initial research efforts in MedVQA adopt models that have proven effective in general VQA domain, and adapt them for medical applications. In terms of image feature extraction, researchers often rely on pre-trained models like VGGNet [1] and ResNet [2], which are fine-tuned for the specific task of MedVQA. On the text side, GRU [3] and LSTM [4] are commonly used to extract textual features, while some approaches incorporate additional semantic information derived from medical corpora to enhance the embeddings used for question representation.
@@ II-B Multi-modal Fusion
In the Visual Question Answering task, multi-modal fusion plays a crucial role by integrating visual and textual features to enable accurate classification. The performance of VQA models largely depends on the effective intra-modal feature extraction and the subsequent inter-modal fusion.
Recent approaches to multi-modal fusion in VQA have introduced various methods to enhance the integration of visual and textual features. A foundational method is SAN[7], which iteratively refines attention on relevant image regions based on the given question. Bilinear pooling has been a major focus in improving fusion by capturing complex interactions between modalities. Yu et al. [8] proposed MFB to address the high computational cost of bilinear pooling by factorizing the interaction into two low-rank matrices, preserving performance while reducing complexity. MUTAN[9] builds on this by applying Tucker decomposition to further compress the bilinear tensor and get a more compact and efficient multi-modal representation. To make full use of bilinear attention maps, Kim et al.[10] proposed BAN, which can capture dependencies between visual and textual features efficiently.
The advent of Transformer[11] has brought a significant shift in multi-modal fusion strategies. Methods like LXMERT[12], hi-VQA[13], MCAN[14] and METER[15] utilize cross-modality Transformers with separate encoders for vision and language, followed by a cross-modality encoder to fuse the extracted features. In the domain of Medical Visual Question Answering, Liu et al.[16] employs a Transformer-based architecture that directly fuse image and text features to generate joint representation. While Transformers are effective for fusion, they require more computational resources and are more complex compared to bilinear pooling-based methods. This makes balancing performance and computational cost particularly important in Medical Visual Question Answering, where medical data is often limited.
@@ III-B Multi-modal Feature Extraction
Question Encoder. In this work, We adopt the pre-trained BioBERT model [19] as our question encoder due to its suitability for processing complex biomedical language. BioBERT, based on the BERT architecture, was fine-tuned on a large and diverse biomedical corpus, and demonstrated strong ability in domain-specific biomedical language representation. Therefore, it can generate more accurate representations of medical questions than general-purpose language models.
@@
While BiomedCLIP also offers a text encoder, we chose not to use it for question encoding in this work. Although the BiomedCLIP Text Encoder aligns image and text features within a shared feature space, it lacks the word-level granularity necessary for capturing detailed linguistic information. This level of detail is critical for the fusion methods introduced in Section III-C, which benefit from precise, word-level representations.
@@ III-C Orthogonal Multi-head Bilinear Attention Network
Our proposed Orthogonal Multi-head Bilinear Attention Network (OMniBAN) integrates a single-layer multi-head self-attention mechanism with bilinear attention networks to efficiently fuse visual and textual features for Medical Visual Question Answering. This design can help capture complex intra-modal and cross-modal interactions effectively while maintaining computational efficiency. By leveraging orthogonal multi-head attention, OMniBAN enhances feature diversity and maximizes information extraction across modalities.
@@ III-C2 Cross-modal Bilinear Attention
where $\circ$ denotes the Hadamard (element-wise) product, allowing for efficient alignment between corresponding elements in the image and question features.
@@
In this formulation, each attention glimpse $h$ learns specialized cross-modal relationships to capture diverse interaction patterns between visual and textual features.
@@ III-C3 Orthogonality Loss
To ensure diverse information captured by the model, we introduce an Orthogonality Loss[20] to encourage each attention glimpse to focus on unique aspects of the input. This loss reduces redundancy across glimpses, which is beneficial considering the subtle and often highly similar pixels present in radiological or pathological medical images. Additionally, it helps address with over-fitting during training, which is common in the field of MedVQA.
In detail, we directly apply Orthogonality Loss ($\mathcal{L}_{\perp}$) to the attention distributions obtained from the bilinear attention, as shown in Algorithm 1. For each pair of attention distributions (glimpses), the inner product of their normalized vectors is computed and squared, with these values summed to form the final orthogonality loss.
@@
10:$\mathcal{L}_{\perp}\leftarrow 0$
@@
14:     Update orthogonality loss: $\mathcal{L}_{\perp}\leftarrow\mathcal{L}_{\perp}+\rho^{2}$
@@ III-C4 Classifier and Loss Function
The joint representation output from BAN is then fed into a classifier to predict the most probable answer. A simple feed-forward neural network is used as the classifier, consisting of two fully connected layers with an intermediate activation function.
@@
The main loss function for this task is Binary Cross-Entropy with Logits Loss, which is a common choice for multi-label classification. The total loss during training combines the main classification loss ($\mathcal{L}_{\bullet}$) with the Orthogonality Loss ($\mathcal{L}_{\perp}$), which encourages both accurate predictions and diversified attention features.
$\mathcal{L}=\mathcal{L}_{\bullet}+\alpha\cdot\mathcal{L}_{\perp}$ (9)
where $\alpha$ is the linear threshold for Orthogonality Loss.
III-C5 Theoretical Complexity Analysis
Let $\mathbf{V}\in\mathbb{R}^{N_{v}\times d_{v}}$ represent the visual features extracted from $N_{v}$ image regions ($N_{v}=1$ in the context of CLIP-based visual backbones), and let $\mathbf{Q}\in\mathbb{R}^{N_{q}\times d_{q}}$ denote the textual features for $N_{q}$ tokens. We analyze and compare the computational complexity of two fusion paradigms: a vanilla Co-Attention mechanism[14][21] (which is implemented as a stack of Transformer layers) and our proposed OMniBAN framework.
Both Transformer-based Co-Attention and OMniBAN networks begin by applying intra-modal self-attention to refine the visual and textual features independently. This initial processing step has a computational complexity of:
$\mathcal{O}_{\text{self-att}}=\mathcal{O}(N_{v}^{2}d_{v}+N_{q}^{2}d_{q})$ (10)
Transformer Co-Attention. A standard Transformer-based co-attention layer processes visual and textual features through cross-modal attention and feed-forward networks (FFN). The dominant operations per layer include:
  • Cross-modal Attention: Projecting features and performing attention between sequences of length $N_{v}$ and $N_{q}$ with dimensions $d_{v}$ and $d_{q}$ results in a cost scaling with:
    $\mathcal{O}_{\text{cross-att}}=\mathcal{O}(N_{q}N_{v}d_{q}+N_{v}d_{v}d_{q}+N_{q}d_{q}^{2})$ (11)
    where $\mathcal{O}(N_{q}N_{v}d_{q})$ denotes the core interaction matrix multiplication cost, assuming that attention dimension is comparable to $d_{q}$.
  • FFN: For a sequence of length $N$ and dimension $d$, this costs $\mathcal{O}(Nd^{2})$. In the co-attention model, this is typically applied to the output sequences of both modalities, and its cost is:
    $\mathcal{O}_{\text{ffn}}=\mathcal{O}(N_{q}d_{q}^{2}+N_{v}d_{v}^{2})$ (12)
For a stack of $L$ co-attention layers, the total fusion complexity can be expressed as:
$\begin{split}\mathcal{O}_{\text{co-att}}&=\mathcal{O}(N_{v}^{2}d_{v}+N_{q}^{2}d_{q})\\ &+L\cdot(\mathcal{O}(N_{q}N_{v}d_{q}+N_{v}d_{v}d_{q}+N_{q}d_{q}^{2}+N_{v}d_{v}^{2}))\end{split}$ (13)
where the dominant terms are:
$\mathcal{O}(N_{v}^{2}d_{v}+N_{q}^{2}d_{q}+LN_{q}N_{v}\max(d_{v},d_{q})+L(N_{q}d_{q}^{2}+N_{v}d_{v}^{2}))$ (14)
OMniBAN Fusion. Following the same intra-modal self-attention, OMniBAN employs Factorized Bilinear Attention Networks [10] across $\gamma$ glimpses. Each glimpse includes:
  • Factorized Bilinear Interaction: Computing cross-modal interactions with a cost much lower than $\mathcal{O}(N_{v}N_{q}d_{v}d_{q})$ through factorization and projection to a lower intermediate dimension $d_{m}$ ($d_{m}\ll\min(d_{v},d_{q})$), which scales approximately with $\mathcal{O}(N_{q}N_{v}d_{m})$.
  • Post-interaction Projection: A simple projection layer after the bilinear step costs $\mathcal{O}(N_{q}d_{q}^{2})$.
Therefore, the total fusion complexity is:
$\mathcal{O}_{\text{OMniBAN}}=\mathcal{O}(N_{v}^{2}d_{v}+N_{q}^{2}d_{q})+\gamma\cdot(\mathcal{O}(N_{q}N_{v}d_{m})+\mathcal{O}(N_{q}d_{q}^{2})))$ (15)
Complexity Comparison. Assuming $d_{v}\approx d_{q}\approx d$, we can further compare the dominant computational costs of Transformer-based co-attention and OMniBAN. The initial intra-modal self-attention, which costs $\mathcal{O}((N_{v}^{2}+N_{q}^{2})d)$, is common to both frameworks and less dominant than the costs of the repeated fusion steps for $L,\gamma>1$. When $L\approx\gamma$, the total complexity difference is primarily determined by the cost per repeated step. OMniBAN’s step is more efficient due to:
  • Cross-modal Interaction: Replacing Transformer’s attention $\mathcal{O}(N_{q}N_{v}d)$ with factorized bilinear $\mathcal{O}(N_{q}N_{v}d_{m})$, where $d_{m}\ll d$.
  • Post-interaction Operation: Using a simpler projection $\mathcal{O}(N_{q}d^{2})$ compared to Transformer’s FFN $\mathcal{O}((N_{q}+N_{v})d^{2})$. While both exhibit the same asymptotic complexity concerning $N_{q}$ and $d_{q}$, the Transformer’s FFN involves a significantly larger constant factor (due to an expanded intermediate layer), which may lead to higher practical FLOPs.
We also empirically confirm these theoretical efficiency gains in Section IV, where we observe that OMniBAN has fewer parameters and reports much smaller FLOPs compared with co-attention.
@@ IV Experiments
IV-A Datasets and Metrics
We conduct our experiments on two public medical VQA datasets: VQA-RAD[22] and SLAKE[23]. VQA-RAD contains 3,515 QA pairs derived from 315 radiology images, and is split into 3,064 QA pairs for training and 451 for testing. SLAKE is a Chinese–English bilingual dataset featuring 642 radiology images and a total of 7,033 QA pairs. It includes richer image modalities and covers a broader range of body parts in its questions. For this study, we focus on the English subset of SLAKE (marked as SLAKE-EN in Table II), which consists of 4,919 QA pairs from 450 images in the training set and 1,061 QA pairs from 96 images in the test set. Both VQA-RAD and SLAKE organize their questions into two types: open-ended and closed-ended. Closed-ended questions typically have a limited set of possible responses (most commonly yes/no), while open-ended questions involve more varied answers.
Since MedVQA can be considered as a multi-label classification task, we adopt a binary cross-entropy (BCE) loss function during training. We primarily use accuracy as our evaluation metric, which is computed as the ratio of correctly predicted answers to the total number of questions. To provide a more comprehensive assessment, we report three separate accuracy scores: overall, open-ended, and closed-ended. This breakdown helps compare model performance across different question types.